OpenStudy (anonymous):

sec x / (1 + sec x) = (1- cos x) / sin ^2 x

OpenStudy (anonymous):

\[\huge \frac{\sec x}{1+\sec x}=\frac{\frac{1}{\cos x}}{1+\frac{1}{cosx }}\]

OpenStudy (anonymous):

\[\huge \frac{1}{(\cos x)\left( 1+\frac{1}{\cos x} \right)}\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[\huge =\frac{1}{1+ \cos \ x}\]

OpenStudy (anonymous):

\[\huge =\frac{1}{1 + \cos \ x}\times\frac{1-\cos \ x}{1- \cos \ x}\]

OpenStudy (anonymous):

Maybe you can do it from here? :>

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

:)